### v1

# Nek Minit: Harnessing Pragmatic Metacognitive Prompting for Explainable Sarcasm Detection of Australian and Indian English

Abstract
Sarcasm is a challenge to sentiment analysis because of the incongruity between stated and implied sentiment. The challenge is exacerbated when the implication may be relevant to a specific country or geographical region. Pragmatic metacognitive prompting (pmp) is a cognition-inspired technique that has been used for pragmatic reasoning. In this paper, we harness pmp for explainable sarcasm detection for Australian and Indian English, alongside a benchmark dataset for standard American English. We manually add sarcasm explanations to an existing sarcasm-labeled dataset for Australian and Indian English called besstie, and compare the performance for explainable sarcasm detection for them with flute, a standard American English dataset containing sarcasm explanations. Our approach utilising pmp when evaluated on two open-weight LLMs (gemma and llama) achieves statistically significant performance improvement across all tasks and datasets when compared with four alternative prompting strategies. We also find that alternative techniques such as agentic prompting mitigate context-related failures by enabling external knowledge retrieval. The focused contribution of our work is utilising pmp in generating sarcasm explanations for varieties of English.
Nek Minit: Harnessing Pragmatic Metacognitive Prompting for Explainable Sarcasm Detection of Australian and Indian English
Ishmanbir Singh* Dipankar Srirag* Aditya Joshi University of New South Wales, Sydney, Australia ishman.singh@student.unsw.edu.au, d.srirag, aditya.joshi@unsw.edu.au Equal contribution
1 Introduction
Sarcasm is a form of verbal irony used to express contempt or ridicule, often by saying the opposite of what one means Joshi et al. (2017). This paper focuses on explainable sarcasm detection as a generation task where, given a text, the LLM must predict if it contains sarcasm, and, if it does, it must generate a textual explanation Kumar et al. (2022). However, sarcasm is socio-culturally situated, and its interpretation often depends on local conventions Oprea and Magdy (2020). This may be evidenced via phrases (such as ‘eshay’ or ‘nek minit111https://en.wikipedia.org/wiki/Nek_minnit; Accessed on 20th May 2025. in the case of Australian English) or statements (such as ‘The sun is out and I am at work yay’ may not be understood as sarcastic by Indian English speakers). Recent work shows that large language models (LLMs) may misinterpret sarcasm in non-Western cultural contexts Atari et al. (2023); Tao et al. (2024). Therefore, we focus on explainable sarcasm detection for two varieties of English: Australian (native variety) and Indian English (non-native variety), alongside standard American English. The extended discussion on the related works is provided in Appendix A.
We harness pragmatic metacognitive prompting (pmp; Lee et al., 2025), an extension of metacognitive prompting Wang and Zhao (2024), to incorporate contextual understanding with respect to the geographical region for the task of explainable sarcasm detection. We compare pmp alongside four prompting baselines for a standard American English dataset (flute; Chakrabarty et al., 2022) and corresponding Australian and Indian English subsets from besstie Srirag et al. (2025a) using two open-weight LLMs. Our proposed approach based on pmp significantly (p0.001) improves the performance on both tasks across all datasets222We provide the prompts used for our experiments in Appendix B.. The novelty of this paper is three-fold: (a) We are the first to explore sarcasm explanation generation in the context of varieties of English; (b) we are the first to explore pmp and kg in the context of sarcasm explanation generation.; (c) we release a manually created dataset of sarcasm explanations for Australian and Indian English. Our findings corroborate past work that shows degradation in performance for language varieties other than standard American English for other tasks Blodgett et al. (2016); Ziems et al. (2023); Joshi et al. (2025); Srirag et al. (2025b).
2 Proposed Approach
Given an input text, our objective is explainable sarcasm detection. This spans two tasks: (a) sarcasm detection, and (b) the generation of a textual explanation when sarcasm is detected. Additionally, the model may output two auxiliary labels: not_sarcastic, if no sarcasm is found; and needs_context, if the model is uncertain and requires additional information to make a judgment. As the texts sampled are all labeled to have sarcasm, the expected output is an explanation for sarcasm for all test samples. Our evaluation methodology for the task is outlined in Figure 1.
Subset Samples flute 150 19.1 28.2 besstie-au 150 47.7 25.7 besstie-in 150 14.3 29.6
2.1 Extending besstie
As shown in Figure 1(a), we annotate 150 sarcastic examples each from Australian and Indian English subsets of besstie with text-based explanations. Two authors of this paper, one native to Australia and one to India, served as annotators to independently write these sarcasm explanations for texts from their corresponding regional subset. Annotators were instructed to provide semi-structured explanations in a format stylistically aligned with flute. This enables consistency in explanation structure across all samples. We refer to these subsets as besstie-au and besstie-in respectively. We also randomly sample 150 sarcastic instances and their corresponding explanations from flute as representative of standard American English. Table 1 reports basic statistics of the evaluation datasets, including the average length (in whitespace-tokenised words) of the sarcastic text and its explanation across the three subsets.
2.2 pmp for Explainable Sarcasm Detection
As shown in Figure 1(b), we structure the prompt as a five-step reasoning scaffold based on pragmatic linguistic theory Lee et al. (2025)333We provide an example output in the Appendix C.. While Lee et al. (2025) utilise pmp for sarcasm detection, we adapt pmp to sarcasm explanation generation for language varieties (Australian English and Indian English) as follows.
1.
Comprehension of Context/Understanding: The LLM identifies key situational elements that may trigger sarcastic meanings, enabling the deduction of intended meanings even for unfamiliar words or expressions.
2.
General Pragmatic Analysis: In this step, the LLM clarifies the true intent in the text by separating genuine beliefs from exaggerated or pretended attitudes. This can further clarify the meaning of unusual or unclear words or expressions.
3.
Preliminary Judgment: An initial hypothesis explanation for sarcasm is generated as a focused seed to guide interpretation.
4.
Meta-Comprehension: The LLM validates the hypothesis explanation by ensuring all relevant contextual cues (both explicit and implicit) are correctly interpreted.
5.
Specific Pragmatic Reassessment: Finally, the LLM systematically examines detailed pragmatic elements, improving and justifying the final explanation, further resolving any residual ambiguity.
[…]
3 Experiment Setup
We experiment with two open-weight decoder-only model optimised for instruction following, namely, Gemma3-12B-Instruct (gemma; Gemma Team et al., 2025) and Llama-3.2-3B-Instruct (llama; Grattafiori et al., 2024)444Text Generation Settings: max_new_tokens=1024, temperature=1.0, top_p=0.95, top_k=64. All experiments are performed using one A100 GPU. We compare pmp with four baseline methods:
1.
zero: We prompt the model directly with a task instruction for sarcasm detection and explanation, without any examples.
2.
few: We include five manually curated examples from flute to the zero prompt. Each example contains a sarcastic sentence paired with an explanation, and these examples are fixed across all evaluations.
3.
origin: We append the geographical origin to the zero prompt to help the model incorporate cultural and linguistic cues specific to the source variety.
4.
kg: We implement an agentic prompting strategy inspired by ReAct Yao et al. (2023). The model is prompted to identify knowledge gaps, integrate information retrieved through search queries, and then respond. The model uses the DuckDuckGo tool555https://github.com/deedy5/duckduckgo_search; Accessed on 19 May 2025 to issue search queries. This method of prompting minimises the gap in context and knowledge from local culture, idioms, or events666An example output is provided in the Appendix D..
We measure performance across three metrics: (a) accuracy (the proportion of correctly predicted labels over all instances); (b) similarity (average cosine similarity between the Sentence-BERT embeddings of the reference explanation and generated explanation; Reimers and Gurevych, 2019); and judge. For judge, we employ GPT-4o OpenAI et al. (2024) with default parameters as an evaluator to assess explanation quality. Given a sarcastic text sample, the corresponding ground truth explanation, and the generated explanation, GPT-4o assigns a score from 0 to 5 based on an arbitrary scoring criterion defined in Appendix E. While accuracy measures the sarcasm detection performance, similarity and judge evaluate the generated explanations.
4 Results
Our results address three questions: (a) How do baseline prompting strategies perform on sarcasm detection and explanation generation?; (b) Does pmp improve over these strategies, particularly for sarcasm explanation task?; and (c) Under what settings does pmp yield the best performance?
Table 2 reports results across the three datasets for the two tasks. For both models, zero and few yield high results on flute for sarcasm detection, with gemma achieving 0.97 and 0.98 accuracy respectively, and llama yielding 0.61 and 0.69. However, these methods degrade substantially on the besstie subsets. Particularly, the performance of gemma when tested on besstie-in remains at 0.59 for both methods, while llama shows a slight degradation (zero: 0.49; few: 0.42). Appending geographic cues in origin prompt yields inconsistent improvements when compared to zero. For example, gemma reports a minor improvement when tested on besstie-au (zero: 0.7; origin: 0.72), llama reports a higher performance gain when tested on besstie-in (zero: 0.49; origin: 0.55). These findings indicate that appending geographical cues alone is insufficient to capture cultural nuances. For explanation generation, the trends among baselines remain consistent. gemma, when prompted with kg, demonstrates improvements in explanation quality on besstie-in (similarity: 0.59; judge: 3.31), suggesting some benefit from agentic reasoning. However, these gains do not consistently transfer across datasets. Moreover, kg is not compatible with llama, limiting its applicability.
[…]
pmp significantly improves over all baselines across both tasks and models. For gemma, pmp achieves the highest detection accuracy (1.00) and explanation quality (judge: 4.92) on flute, and similarly strong performance on besstie-au (accuracy: 0.94; judge: 3.98) and besstie-in (accuracy: 0.91; judge: 3.76). For llama, pmp significantly improves over baseline performances: on besstie-in, it improves similarity and judge to 0.52 and 2.47 compared to origin(0.41 and 1.69), and accuracy of 0.79 from 0.55. We also note that while each row of the table reports lower values for Australian and Indian English as compared to standard American English present in flute, our results on pmp method highlight the potential to bridge the performance gap between varieties. The improvements across all datasets are statistically significant (p0.001) relative to the zero baseline.
Table 3 presents an error analysis of two types of explanation errors: instances incorrectly classified as ns, and those requiring nc. On both besstie datasets, pmp substantially reduces these errors. On besstie-in, the count of ns cases drops from 18 under zero to 1 with pmp; similarly, nc cases reduce from 43 to 13. These results are consistent for besstie-au, confirming that pmp is particularly effective in cases requiring nuanced cultural or pragmatic interpretation. However, for besstie-in, models more often output nc, especially for zero and few. This indicates that the model often cannot interpret the sarcasm from surface-level language alone, likely due to cultural or regional differences in how sarcasm is expressed. Here, kg proves effective by supplying external context that helps the model bridge this gap, reducing the ns errors to 4, when compared to 43 from zero.
5 Conclusion
We demonstrated how pragmatic metacognitive prompting (pmp) can be harnessed to generate sarcasm explanations for text written in varieties. We evaluate our approach on three language varieties of English: standard American English, Australian and Indian English, and two open-source LLMs: gemma and llama. We annotated an existing dataset for the latter two with sarcasm explanations to assess model performance for explainable sarcasm detection. Standard prompting methods, such as zero-shot and few-shot, perform well on flute, but failed to generalise to Australian and Indian English subsets of besstie. pmp significantly (p0.001) improved the performance across both tasks, both Australian and Indian English, and both the models. Agentic prompting methods like kg also reduced context-related failures by enabling dynamic knowledge integration. Our results and error analysis demonstrated the limitations of generic prompts and the importance of pragmatic scaffolding for figurative language understanding. Our findings suggest that reasoning-aware prompting offers a viable pathway to improve explanation generation for sarcasm.
Limitations
We perform sarcasm detection only on positive text samples, i.e. sarcastic text, as this study primarily explores sarcasm explanation as a task. There may be multiple possible explanations for sarcasm. Also, we acknowledge that there are varieties within a country related to the state, native language, and so on. Note that we perform sarcasm detection only on positive text samples, i.e. sarcastic text, as this study primarily explores sarcasm explanation as a task. However, as a first study on sarcasm explanation for language varieties, our observations are a good starting point for the future. The proposed pmp approach helps besstie-au and besstie-in to retrieve appropriate context which is not necessary in the case of flute based on past findings in the bias of LLMs towards a western-centric context.
Ethical Considerations
The research was approved by the ethics board of the host institution. All annotators were native speakers of the respective English varieties (Australian and Indian) and participated voluntarily in the annotation process. Given that sarcasm is often context-dependent and culturally embedded, we acknowledge the sensitivity involved in interpreting or misclassifying user-generated content. No personally identifiable information was used in this study.
Acknowledgment
This paper is the outcome of a Taste-of-Research scholarship awarded to Ishmanbir Singh by the Faculty of Engineering at UNSW Sydney.
Appendix A Related Work
Sarcasm detection has employed statistical models Joshi et al. (2017), sentiment-incongruity heuristics Riloff et al. (2013), and traditional neural architectures such as LSTMs Ghosh et al. (2017), CNNs Poria et al. (2016), GNNs Lou et al. (2021); Liang et al. (2022) and Transformer Yao et al. (2024). Prompting LLMs has shown promise in guiding models to reason about sarcasm effectively Liu et al. (2023); Yao et al. (2025). Specifically, Pragmatic Metacognitive Prompting (pmp) Lee et al. (2025) introduces a structured approach that mirrors human pragmatic reasoning by incorporating reflection and analysis of implied meanings, contextual cues, and speaker intent. This method has demonstrated improved performance in sarcasm detection tasks but has not been used for sarcasm explanation generation. Additionally, agentic prompting methods like ReAct Yao et al. (2023) enable models to actively retrieve and integrate external knowledge, facilitating context-aware reasoning in sarcasm detection. Beyond sarcasm detection, sarcasm explanation generation has been investigated using datasets like flute Chakrabarty et al. (2022) which provide figurative language instances with explanations. Models such as TEAM Jing et al. (2023) employ multi-source semantic graphs to generate multimodal sarcasm explanations, integrating visual and textual cues. Ours is the first work that examines the different prompting strategies for the explanation generation of sarcasm for language varieties.
Appendix B Prompts
In this paper, we evaluate models on the task of explainable sarcasm detection, by prompting models to first identify sarcasm and then provide an explanation for classifying the text to be sarcastic. Below are the prompts that we used.
B.1 zero
For the provided text, perform one of the tasks. If the text is not sarcastic, output ‘not_sarcastic’. If the text is sarcastic, provide an explanation in one or two sentences. Output ‘need_context’ if you cannot explain the sarcasm.
B.2 few
For the provided text, perform one of the tasks. If the text is not sarcastic, output ‘not_sarcastic’. If the text is sarcastic, provide an explanation in one or two sentences. Output ‘need_context’ if you cannot explain the sarcasm. Following are some examples of sarcastic texts and their explanations.
1.
Text: I got the greatest news today when I found out my company is getting rid of me after 10 years of faithful service
Explanation: sarcastic. Being fired from a job after 10 years of service is not something to be happy about and hence receiving the greatest news is sarcastic. -
2.
Text: I feel great that I am single and have not accomplished anything significant inspite of being 30 years old
Explanation: sarcastic. Someone who is 30 years old and single has not achieved anything significant in their life so far and so should not feel honored by that fact. -
3.
Text: The idea of the genius Republican party blatantly rejecting climate change and its effects on our planet makes me feel safe for the future
Explanation: sarcastic. The Republican party’s denial of climate change will have catastrophic effects on our planet in the future and so to feel confident about the future is absurd. -
4.
Text: I was so impressed when my hair dresser tried a new product on my hair but the end result was me getting to shave all my hair off
Explanation: sarcastic. Shaving all your hair off is not an impressive hairstyle and is often seen as a last resort when someone has messed up your hair. -
5.
Text: I am on seventh heaven because I don’t have enough money to pay my important bill.
Explanation: sarcastic. Having not enough money to pay an important bill creates a lot of stress and anxiety and is definitely not a heavenly feeling.
B.3 origin
This text is from [Australian/Indian] subreddit post or comment. If the text is not sarcastic, output ‘not_sarcastic’. If the text is sarcastic, provide an explanation in one or two sentences. Output ‘need_context’ if you cannot explain the sarcasm.
B.4 kg
Think step-by-step and feel free to use tools whenever they would be helpful. Once you have finished reasoning and using tools:
If the text is not sarcastic:
{‘‘action’’: ‘‘Final Answer’’, ‘‘action_input’’: ‘‘not_sarcastic’’} -
If the text is sarcastic:
{‘‘action’’: ‘‘Final Answer’’, ‘‘action_input’’: ‘‘sarcastic. Explanation: concise explanation (1-2 sentences)’’} -
If you cannot explain the sarcasm:
{‘‘action’’: ‘‘Final Answer’’, ‘‘action_input’’: ‘‘need_context’’}
All outputs must be valid JSON, with no unescaped double quotes; use single quotes. You have access to the following tools:
Search: Tool for getting up to date answers to current or historical events and word/phrase definitions., args: {‘tool_input’: {‘type’: ‘string’}}
Use a json blob to specify a tool by providing an action key (tool name) and an action_input key (tool input). Valid ‘‘action’’ values are: ‘‘Final Answer’’ or ‘‘Search’’.
Provide only ONE action per $JSON_BLOB, as shown:
{‘‘action’’: $TOOL_NAME, ‘‘action_input’’: $INPUT}
Follow this format:
Question: input question to answer
Thought: consider previous and subsequent steps
Action:
$JSON_BLOB
Observation: action result
... (repeat Thought/Action/Observation N times)
Thought: I know what to respond
Action:
{‘‘action’’: ‘‘Final Answer’’, ‘‘action_input’’: ‘‘Final response to human’’}
Begin! Reminder to ALWAYS respond with a valid json blob of a single action. Use tools if necessary. Respond directly if appropriate. Format is
Action:$JSON_BLOB then Observation:.
Thought:
Human:
B.5 pmp
For the provided text, perform one of the tasks. If the text is not sarcastic, output ‘not_sarcastic’. If the text is sarcastic, follow these steps in your analysis:
1.
Comprehension of Context/Understanding:
Repeat the text and summarize it to confirm your understanding
Identify the context, topic, and key elements of the situation described
2.
General Pragmatic Analysis:
What does the author imply about the situation with their statement?
What does the author think about the situation?
Are what the author implies and what they think saying the same thing?
Is the author pretending to have a certain attitude toward the conversation?
3.
Preliminary Judgment:
Based on your analysis, formulate an initial explanation of why the text is sarcastic
4.
Comprehension of Preliminary Judgment/Context:
Summarize your preliminary analysis
Ensure you’ve captured all relevant contextual elements
5.
Specific Pragmatic Analysis/Reassessment:
Implicature: What is implied beyond the literal meaning?
Presuppositions: What information is taken for granted?
Intent: What does the author hope to achieve with their statement?
Polarity: Is there a mismatch between apparent and intended tone?
Pretense: Is there pretense in the author’s attitude?
Meaning gap: What is the difference between literal and implied meaning?
Context: Consider slang, cultural references, or communication patterns
6.
After completing your analysis, perform one of the tasks.
If the text is sarcastic, provide an explanation in one or two sentences.
Output ‘need_context’ if you cannot explain the sarcasm.